OpenStudy (anonymous):

Find the angle between a and b. Given that a= 6i - 2j -3k and b= i+j+k

OpenStudy (anonymous):

3d hvn't learnt yet

OpenStudy (anonymous):

arccos( a dot b) / ||a|| * ||b|| arcos(6*1 - 2*1 + 3*1/sqrt(36+4+9)*sqrt(3)) = arccos(7/(7*sqrt(3))) = arccos(1/sqrt(3)) = 54.7

myininaya (myininaya):

\[\cos(\theta)=\frac{a \cdot b}{|a| |b|} => \theta=\cos^{-1}(\frac{a \cdot b}{|a||b|})\]

myininaya (myininaya):

so we need to find a dot b \[a \cdot b=6(1)+(-2)(1)+(-3)(1)=-7\]

OpenStudy (amistre64):

is equal to or greater than theta?

myininaya (myininaya):

we need to find |a| \[|a|=\sqrt{6^2+(-2)^2+(-3)^2}=\sqrt{36+4+9}=\sqrt{49}=7\] \[|b|=\sqrt{1^2+1^2+1^2}=\sqrt{3}\]

myininaya (myininaya):

\[\theta=\cos^{-1}( \frac{-7}{7 \sqrt{3}})=\cos^{-1}(\frac{-1}{\sqrt{3}})=\cos^{-1}(-\frac{\sqrt{3}}{3})\]

myininaya (myininaya):

this is the angle created by the vectors a and b so = unless you use an approximation and depending on your approximation theta could be greater than or less than your approximation

myininaya (myininaya):



OpenStudy (amistre64):

curvey vector lol

myininaya (myininaya):

lol

myininaya (myininaya):

i guess i could have use that draw a straight line tool